# Scaling Laws for Task-Optimized Models of the Primate Visual Ventral Stream

### v2

@@ Abstract
When trained on large-scale object classification datasets, certain artificial neural network models begin to approximate core object recognition (COR) behaviors and neural response patterns in the primate visual ventral stream (VVS). While recent machine learning advances suggest that scaling model size, dataset size, and compute resources improve task performance, the impact of scaling on brain alignment remains unclear. In this study, we explore scaling laws for modeling the primate VVS by systematically evaluating over 600 models trained under controlled conditions on benchmarks spanning V1, V2, V4, IT and COR behaviors. We observe that while behavioral alignment continues to scale with larger models, neural alignment saturates. This observation remains true across model architectures and training datasets, even though models with stronger inductive bias and datasets with higher-quality images are more compute-efficient. Increased scaling is especially beneficial for higher-level visual areas, where small models trained on few samples exhibit only poor alignment. Finally, we develop a scaling recipe, indicating that a greater proportion of compute should be allocated to data samples over model size. Our results suggest that while scaling alone might suffice for alignment with human core object recognition behavior, it will not yield improved models of the brain’s visual ventral stream with current architectures and datasets, highlighting the need for novel strategies in building brain-like models.  [^1] We open-source all code and checkpoints: github.com/epflneuroailab/scaling-primate-vvs.
1 Introduction
@@
The advent of neural networks has revolutionized our understanding and modeling of complex neural processes. A particularly active area of study is the ventral visual stream in primates, a key pathway in the brain responsible for processing visual information Goodale & Milner (1992); Grill-Spector et al. (2001); Malach et al. (2002); Kriegeskorte et al. (2008). Neural networks, when trained on extensive datasets, have emerged as the most accurate quantitative tools for simulating the response patterns of neurons within this stream Yamins et al. (2014); Schrimpf et al. (2018). These advanced models offer a precise computational account of how neural mechanisms in the brain give rise to visual perception.
@@
Recent developments in machine learning have emphasized the significance of both the volume of training data and the complexity of model architectures Kaplan et al. (2020); Hoffmann et al. (2022); Zhai et al. (2022); Bahri et al. (2022); Antonello et al. (2023); Muennighoff et al. (2023); Aghajanyan et al. (2023); Isik et al. (2024). Henighan et al. (2020). These findings raise the question: Can we build better models of the brain by scaling up model architectures and dataset sizes? Recent studies have found that in pre-trained models, the number of parameters and dataset samples respectively seem to improve predictions of fMRI and behavioral measurements (Antonello et al., 2023; Muttenthaler et al., 2023). With the numerous differences between pre-trained models however, the relative contributions of model parameters and dataset size to brain and behavioral alignment are not clear.
@@
In this paper, we examine how scaling – of model parameters and training dataset size – impacts the alignment of artificial neural networks with the primate ventral visual stream. We systematically train models from a variety of architectural families on image classification datasets which allows us to independently control and observe the effects of model complexity and data volume. To capture the observed trends, we introduce parametric power-law trends that describe the impact of scale on alignment with behavior and brain regions along the visual ventral stream.
We summarize the contributions of this work as follows:
@@
  • While scale initially improves alignment, brain alignment saturates whereas behavioral alignment continues to improve.
@@
  • Architectures with stronger inductive bias (such as convolutions) and datasets with higher-quality images are more sample- and compute-efficient.
@@
Figure 1: a) For a given compute budget ($C$), we determine the scaling laws for maximal neural and behavioral alignment to the primate visual ventral stream. b) We find consistent scaling laws for brain and behavioral alignment across over 600 models. While we predict models to approach perfect behavioral alignment of 1 at large scales, the effect of scaling on brain alignment is already saturating at 0.48.
@@ 2 Related Work
The ventral visual stream, a critical pathway in the primate brain, including humans, plays a key role in visual perception, extending from the occipital to the temporal lobes and serving as the ”what pathway” for object recognition and form representation Goodale & Milner (1992); Grill-Spector et al. (2001); Malach et al. (2002); Kriegeskorte et al. (2008). Beginning in the primary visual cortex (V1), where basic visual information from retinal ganglion cells is processed, the ventral stream proceeds through areas such as V2, V3, V4, and the inferotemporal cortex (IT), each responsible for increasingly complex features of visual perception Kandel et al. (2000).
Despite decades of research and a wealth of brain data, the precise neural mechanisms underlying visual perception are not well understood.
@@
Particular artificial neural networks (ANNs) are the most accurate models of brain responses in the visual ventral stream and associated core object recognition behaviors Schrimpf et al. (2018; 2020). Models optimized for ecologically viable tasks (Yamins & DiCarlo, 2016) in particular have demonstrated strong brain and behavioral alignment (Yamins et al., 2014; Khaligh-Razavi & Kriegeskorte, 2014; Cadena et al., 2019; Schrimpf et al., 2018; Nayebi et al., 2018; Kietzmann et al., 2019; Rajalingham et al., 2018; Zhuang et al., 2021; Geiger et al., 2022) – notably these models are trained purely on image classification datasets, without fitting to brain data.
@@
Recent advancements in artificial intelligence are driven by scaling the model size and training data. Empirical evidence suggests a power-law relationship between model performance and both model parameters and dataset size, indicating that continued scaling will further improve performance Kaplan et al. (2020); Cherti et al. (2023); Zhai et al. (2022); Hoffmann et al. (2022); Dehghani et al. (2023); Henighan et al. (2020); Brown et al. (2020); Bahri et al. (2022); Hestness et al. (2017). The power-law exponents enable the optimal allocation of compute between model parameters and dataset samples, such that performance is maximized Kaplan et al. (2020); Hoffmann et al. (2022).
While scaling laws for machine learning performance has been extensively studied, the scaling laws for brain alignment remain unclear. Recent studies indeed suggest an involvement of both model size and data volume in the functional alignment with brain data Azabou et al. (2023); Benchetrit et al. (2023); Caro et al. (2024); Antonello et al. (2023). Conversely, Muttenthaler et al. Muttenthaler et al. (2023) indicate that sample size is critical for behavioral alignment. We here unify these results, in the realm of the primate visual ventral stream, into quantitative scaling laws for how model and dataset sizes relate to alignment with the brain and behavior.
@@ 3 Methods
To evaluate the alignment of our model with brain function, we utilize a range of benchmarks from Brain-Score Schrimpf et al. (2018; 2020). These benchmarks assess model performance by comparing model activations or behavior with primate neural data using the same images. Specifically, the V1 and V2 benchmarks compare model outputs to primate single-unit recordings from Freeman et al. (2013), using 315 texture images and data from 102 V1 and 103 V2 neurons. For the V4 and IT benchmarks, 2,560 images are used to match model activations to primate Utah array recordings from Majaj et al. (2015), based on data from 88 V4 and 168 IT electrodes. A linear regression is trained on 90% of the images to correlate model and neural data, with prediction accuracy for the remaining 10% evaluated using Pearson correlation, repeated ten times for cross-validation. The behavioral benchmark assesses model predictions for 240 images against primate behavioral data from Rajalingham et al. (2018) using a logistic classifier trained on 2,160 labeled images. Pearson correlation is used to measure the similarity in confusion patterns between model predictions and primate responses. All benchmark scores are normalized to their respective maximum possible values.
We define the model’s alignment score $S$ (and an inverse Misalignment Score $L=1-S$) as the average across the V1, V2, V4, IT, and behavioral benchmark scores. Layers are committed to brain regions based on models trained on a full dataset, and applied to all variants trained with subsampled datasets. As we reused the same neural and behavioral data both to select the optimal model layer for readout and to assess the model’s alignment, we validated the benchmark results on a private split of each dataset on Brain-Score. We observed an almost perfect correlation between the results on the private and public splits (Appendix C).
@@
For our experiments, we selected two image classification datasets: ImageNet Deng et al. (2009) and EcoSet Mehrer et al. (2021). ImageNet, with millions of labeled images across 1,000 categories, has long been a benchmark in computer vision, designed to challenge and evaluate automated visual object recognition systems. On the other hand, EcoSet is a more recent dataset, designed to provide an ecologically valid representation of human-relevant objects. It contains over 1.5 million images spanning 565 basic-level categories, curated to better reflect the natural distribution of objects in the real world, aligning with human perceptual and cognitive experiences.
To create subsets of ImageNet and EcoSet, we sampled $d\in{1,3,10,30,100,300}$ images per category. For $d\in{1,10,100}$, we repeated the runs with three random seeds to ensure robustness. For ConvNeXts (Liu et al., 2022b) and ViTs (Touvron et al., 2022), we used the training recipes developed by the original model authors. The remaining models were trained for 100 epochs using a minibatch size of 512. We employed a stochastic gradient descent (SGD) optimizer with a cosine decaying learning rate schedule, starting with a peak learning rate of 0.1 and incorporating a linear warm-up phase spanning five epochs. We maintained the momentum at 0.9 and applied a weight decay of $10^{-4}$. Cross-entropy loss was used as the minimization objective. We utilized standard ImageNet data augmentations, specifically random resized cropping and horizontal flipping.
@@
Following previous work on scaling laws Zhai et al. (2022); Hoffmann et al. (2022); Besiroglu et al. (2024), we fit power law functions in the form
@@
on the data where $L$ is the misalignment score, and $X$ is an independent variable, such as the number of samples seen ($D$), number of parameters ($N$), and the total training floating point operations (FLOPs) ($C$). Coefficients $E$, $A$, and $\alpha$ are found by minimizing
@@
$(N^{*},\;D^{*})=\operatorname*{arg\,min}_{N,\;D}L(N,\;D)\quad\textrm{s.t}\quad\textrm{FLOPs}(N,\;D)=C$ (4)
In that regard, following Hoffmann et al. (2022); Besiroglu et al. (2024) we fit a parametric function of the form
@@
where the loss ($\hat{L}$) is a function of parameter count ($N$) and number of samples seen ($D$). In Eq. 5, the first term represents the loss in an ideal data generation scenario (entropy), the second and the third terms reflect the under-performance of a model due to limitations in parameter and data size Hoffmann et al. (2022); Muennighoff et al. (2023). Following the example of Hoffmann et al. (2022), we learn variables $\{E,\;A,\;\alpha,\;B,\;\beta\}$ that characterizes misalignment by solving
$\operatorname*{arg\,min}_{e,\;a,\;\alpha,\;b,\;\beta}\quad\sum_{i\in[\#\textrm{Runs}]}\textrm{Huber}_{\delta}\left(\textrm{LSE}(a-\alpha\log N_{i},b-\beta\log D_{i},e)-\log L_{i}\right)$ (6)
@@
Both Kaplan et al. (2020); Hoffmann et al. (2022) assume that compute follows the relationship $C(N,D)\approx 6ND$ to predict the optimal allocation of compute ($C$) to $N$ and $D$ using a set of equations with the learned variables mentioned above:
$\begin{split}&N^{*}(C)=G(C/6)^{a},\quad\quad D^{*}(C)=G^{-1}(C/6)^{b}\\ &\textrm{where}\quad\quad a^{\prime}=\frac{\beta}{\alpha+\beta}\>,\quad b^{\prime}=\frac{\alpha}{\alpha+\beta}\>,\quad G=\left(\frac{\alpha A}{\beta B}\right)^{\frac{1}{\alpha+\beta}}\end{split}$ (7)
However, we observe that $C(N,D)\approx 6ND$ does not hold with different architectures, and various CNN families have a slightly different relationship of $C$, $N$, and $D$. As such, we assume a power-law relationship of the form
@@
$N^{*}(C)=G(C/m)^{a^{\prime}/n},\quad\quad D^{*}(C)=G^{-1}(C/m)^{b^{\prime}/n}$ (9)
@@
Finally, to avoid large constants during curve fitting, we rescale the variables $C$, $N$, and $D$ by setting $\tilde{C}=C/10^{13}$, $\tilde{N}=N/10^{5}$, and $\tilde{D}=D/10^{4}$.
@@ 4.1 Scaling drives behavioral alignment, but saturates for neural alignment
4.2 Architectural Inductive Bias Influences Alignment and Scaling Behavior

### v3

@@ Abstract
When trained on large-scale object classification datasets, certain artificial neural network models begin to approximate core object recognition behaviors and neural response patterns in the primate brain. While recent machine learning advances suggest that scaling compute, model size, and dataset size improves task performance, the impact of scaling on brain alignment remains unclear. In this study, we explore scaling laws for modeling the primate visual ventral stream by systematically evaluating over 600 models trained under controlled conditions on benchmarks spanning V1, V2, V4, IT and behavior. We find that while behavioral alignment continues to scale with larger models, neural alignment saturates. This observation remains true across model architectures and training datasets, even though models with stronger inductive biases and datasets with higher-quality images are more compute-efficient. Increased scaling is especially beneficial for higher-level visual areas, where small models trained on few samples exhibit only poor alignment. Our results suggest that while scaling current architectures and datasets might suffice for alignment with human core object recognition behavior, it will not yield improved models of the brain’s visual ventral stream, highlighting the need for novel strategies in building brain models.
Machine Learning, ICML, NeuroAI, scaling laws, computational neuroscience, computer vision
Figure 1: a) For a given compute budget ($C$), we determine the scaling laws for maximal neural and behavioral alignment to the primate visual ventral stream. b) We find consistent scaling laws for brain and behavioral alignment across over 600 models. While we predict models to approach perfect behavioral alignment at large scales, the effect of scaling on brain alignment is already saturating.
@@
The advent of neural networks has revolutionized our understanding and modeling of complex neural processes. A particularly active area of study is the ventral visual stream in primates, a key pathway in the brain responsible for processing visual information [Goodale1992, GrillSpector2001, Malach2002, Kriegeskorte2008]. Neural networks, when trained on extensive datasets, have emerged as the most accurate quantitative tools for simulating the response patterns of neurons within this stream [Yamins_Hong_Cadieu_Solomon_Seibert_DiCarlo_2014, SchrimpfKubilius2018BrainScore]. These advanced models offer a precise computational account of how neural mechanisms in the brain give rise to visual perception.
@@
Recent developments in machine learning have emphasized the significance of both the volume of training data and the complexity of model architectures [kaplan2020scalingNlp, hoffmann2022Chinchilla, zhai2022scalingVit, bahri2022explainingscaling, antonello2023scalingfmri, muennighoff2023scalingConstrained, aghajanyan2023generative, isik2024Downstream]. These findings raise the question: Can we build better models of the brain by scaling up model architectures and dataset sizes? Recent studies have found that in pre-trained models, the number of parameters and dataset samples respectively seem to improve predictions of fMRI and behavioral measurements [antonello2023scalingfmri, muttenthaler2023human]. With the numerous differences between pre-trained models however, the relative contributions of model parameters and dataset size to brain and behavioral alignment are not clear.
Despite recent successes in using neural networks as models of the brain, a comprehensive understanding of how model scale—separately and jointly across parameters, dataset size, and compute—affects functional alignment with different cortical areas remains elusive. Previous studies have often relied on heterogeneously trained models using off-the-shelf checkpoints [conwell2024], or focused narrowly on specific brain areas (e.g., IT only [linsley2023worseit]), frequently using proxy quantities such as task performance. Our work addresses these limitations through a systematic, from-scratch training protocol spanning over 600 models, enabling controlled comparisons and robust parametric estimation of scaling laws for both behavioral and neural alignment across the entire ventral visual hierarchy. This approach offers a clearer disentanglement of the respective contributions of architecture, data, and optimization objective to brain modeling.
@@
In this paper, we examine how scaling – of model parameters and training dataset size – impacts the alignment of artificial neural networks with the primate ventral visual stream. We systematically train models from a variety of architectural families on image classification datasets which allows us to independently control and observe the effects of model complexity and data volume. To capture the observed trends, we introduce parametric power-law trends that describe the impact of scale on alignment with behavior and brain regions along the visual ventral stream. We summarize the contributions of this work as follows:
@@
  • While scale initially improves alignment, brain alignment saturates. Behavioral alignment on the other hand continues to improve.
@@
  • Architectures with stronger inductive bias (e.g., convolutions and recurrence) and datasets with higher-quality images are more sample- and compute-efficient.
@@ 1 Related Work
The ventral visual stream, a critical pathway in the primate brain, including humans, plays a key role in visual perception, extending from the occipital to the temporal lobes and serving as the ”what pathway” for object recognition and form representation [Goodale1992, GrillSpector2001, Malach2002, Kriegeskorte2008]. Beginning in the primary visual cortex (V1), where basic visual information from retinal ganglion cells is processed, the ventral stream proceeds through areas such as V2, V3, V4, and the inferotemporal cortex (IT), each responsible for increasingly complex features of visual perception [Kandel2000]. Despite decades of research and a wealth of brain data, the precise neural mechanisms underlying visual perception are not well understood.
@@
Particular artificial neural networks (ANNs) are the most accurate models of brain responses in the visual ventral stream and associated core object recognition behaviors [SchrimpfKubilius2018BrainScore, Schrimpf2020integrative]. Models optimized for ecologically viable tasks [Yamins2016] in particular have demonstrated strong brain and behavioral alignment [Yamins_Hong_Cadieu_Solomon_Seibert_DiCarlo_2014, KhalighRazavi2014, Cadena2019, SchrimpfKubilius2018BrainScore, Nayebi2018, Kietzmann2019, rajalingham2018largescale, zhuang2021unsupervised, geigerschrimpf2022wiringup] – notably these models are trained purely on image classification datasets, without fitting to brain data.
@@
Recent advancements in artificial intelligence are driven by scaling the model size and training data. Empirical evidence suggests a power-law relationship between model performance and both model parameters and dataset size, indicating that continued scaling will further improve performance [kaplan2020scalingNlp, cherti2022reproducibleContrastive, zhai2022scalingVit, hoffmann2022Chinchilla, dehghani2023vit22b, henighan2020scalingGenerative, brown2020scalingfewshotnlp, bahri2022explainingscaling, hestness2017deepscaling]. The power-law exponents enable the optimal allocation of compute between model parameters and dataset samples, such that performance is maximized [kaplan2020scalingNlp, hoffmann2022Chinchilla].
While scaling laws for machine learning performance has been extensively studied, the scaling laws for brain alignment remain unclear. Recent studies suggest an involvement of both model size and data volume in the functional alignment with brain data [azabou2023unifiedperceivorio, meta2023brain, caro2024brainlm, antonello2023scalingfmri]. Conversely, muttenthaler2023human indicate that sample size is critical for behavioral alignment. We here unify these results, in the realm of the primate visual ventral stream, into quantitative scaling laws for how model and dataset sizes relate to alignment with the brain and behavior.
@@ 2 Methods
To evaluate the alignment of our model with brain function, we utilize a range of benchmarks from Brain-Score [SchrimpfKubilius2018BrainScore, Schrimpf2020integrative]. These benchmarks assess model performance by comparing model activations or behavior with primate neural data using the same images. Specifically, the V1 and V2 benchmarks compare model outputs to primate single-unit recordings from [freeman2013V1V2], using 315 texture images and data from 102 V1 and 103 V2 neurons. For the V4 and IT benchmarks, 2,560 images are used to match model activations to primate Utah array recordings from [majaj2015V4IT], based on data from 88 V4 and 168 IT electrodes. A linear regression is trained on 90% of the images to correlate model and neural data, with prediction accuracy for the remaining 10% evaluated using Pearson correlation, repeated ten times for cross-validation. The behavioral benchmark assesses model predictions for 240 images against primate behavioral data from [rajalingham2018largescale] using a logistic classifier trained on 2,160 labeled images. Pearson correlation is used to measure the similarity in confusion patterns between model predictions and primate responses. All benchmark scores are normalized to their respective maximum possible values.
We define the model’s alignment score $S$ (and an inverse Misalignment Score L=1SL=1-S) as the average across the V1, V2, V4, IT, and behavioral benchmark scores. Layers are committed to brain regions based on models trained on a full dataset, and applied to all variants trained with subsampled datasets. As we reused the same neural and behavioral data both to select the optimal model layer for readout and to assess the model’s alignment, we validated the benchmark results on a private split of each dataset on Brain-Score. We observed an almost perfect correlation between the results on the private and public splits (Appendix C).
@@
For our experiments, we selected two image classification datasets: ImageNet [deng2009imagenet] and EcoSet [mehrer20122ecoset]. ImageNet, with millions of labeled images across 1,000 categories, has long been a benchmark in computer vision, designed to challenge and evaluate automated visual object recognition systems. On the other hand, EcoSet is a more recent dataset, designed to provide an ecologically valid representation of human-relevant objects. It contains over 1.5 million images spanning 565 basic-level categories, curated to better reflect the natural distribution of objects in the real world, aligning with human perceptual and cognitive experiences.
To create subsets of ImageNet and EcoSet, we sampled $d\in{1,3,10,30,100,300}$ images per category. For $d\in{1,10,100}$, we repeated the runs with three random seeds to ensure robustness. For ConvNeXts [liu2022convnext] and ViTs [deit2022revenge], we used the training recipes developed by the original model authors. The remaining models were trained for 100 epochs using a minibatch size of 512. We employed a stochastic gradient descent (SGD) optimizer with a cosine decaying learning rate schedule, starting with a peak learning rate of 0.1 and incorporating a linear warm-up phase spanning five epochs. We maintained the momentum at 0.9 and applied a weight decay of $10^{-4}$. Cross-entropy loss was used as the minimization objective. We utilized standard ImageNet data augmentations, specifically random resized cropping and horizontal flipping.
@@
Following previous work on scaling laws [zhai2022scalingVit, hoffmann2022Chinchilla, chinchilla2024besiroglu], we fit power law functions in the form
@@
on the data where $L$ is the misalignment score, and $X$ is an independent variable, such as the number of samples seen ($D$), number of parameters ($N$), and the total training floating point operations (FLOPs) ($C$). Coefficients $E$, AA, and $\alpha$ are found by minimizing
@@
$(N^{*},D^{*})=\operatorname*{arg\,min}_{N,D}L(N,D),\>\textrm{FLOPs}(ND)=C$ (4)
In that regard, following [hoffmann2022Chinchilla, chinchilla2024besiroglu] we fit a parametric function of the form
@@
where the loss ($\hat{L}$) is a function of parameter count ($N$) and number of samples seen ($D$). In Eq. 5, the first term represents the loss in an ideal data generation scenario (entropy), the second and the third terms reflect the under-performance of a model due to limitations in parameter and data size [hoffmann2022Chinchilla, muennighoff2023scalingConstrained]. Following the example of hoffmann2022Chinchilla, we learn variables $\{E,\;A,\;\alpha,\;B,\;\beta\}$ that characterizes misalignment by solving
$\operatorname*{arg\,min}_{e,\;a,\;\alpha,\;b,\;\beta}\sum_{i\in[\#\textrm{Runs}]}\textrm{Huber}_{\delta}\Big(\log L_{i}$- (6)
$\textrm{LSE}\big(a-\alpha\log N_{i},b-\beta\log D_{i},e\big)\Big)$
@@
Both kaplan2020scalingNlp, hoffmann2022Chinchilla assume that compute follows the relationship $C(N,D)\approx 6ND$ to predict the optimal allocation of compute ($C$) to $N$ and $D$ using a set of equations with the learned variables mentioned above:
$\begin{split}&N^{*}(C)=G(C/6)^{a},\quad D^{*}(C)=G^{-1}(C/6)^{b}\\ &\textrm{where}\>a^{\prime}=\frac{\beta}{\alpha+\beta},\>b^{\prime}=\frac{\alpha}{\alpha+\beta},\>G=\left(\frac{\alpha A}{\beta B}\right)^{\frac{1}{\alpha+\beta}}\end{split}$ (7)
However, we observe that $C(N,D)\approx 6ND$ does not hold with different architectures, and various CNN families have a slightly different relationship of $C$, NN, and $D$. As such, we assume a power-law relationship of the form
@@
$N^{*}(C)=G(C/m)^{a^{\prime}/n},\>D^{*}(C)=G^{-1}(C/m)^{b^{\prime}/n}$ (9)
@@
Finally, to avoid large constants during curve fitting, we rescale the variables $C$, NN, and $D$ by setting $\tilde{C}=C/10^{13}$, $\tilde{N}=N/10^{5}$, and $\tilde{D}=D/10^{4}$.
@@ 3.1 Scaling drives behavioral alignment, but saturates for neural alignment
3.2 Architectural Inductive Bias Influences Alignment and Scaling Dynamics
@@
Fig. 3b highlights that architectural priors critically shape alignment dynamics, particularly in low-data settings. CORnet models, which incorporate recurrence, achieve relatively high alignment early in training—outperforming both convolutional and transformer-based models under limited supervision. Yet as training data increases, this initial advantage wanes, and alignment scores across architectures begin to converge. This suggests that while certain inductive biases offer sample efficiency, their long-term benefits may be outpaced by deeper or more flexible architectures given sufficient data. Overall, these findings emphasize that strong inductive biases—such as convolution and recurrence—facilitate better alignment when data is limited, whereas extensive task-driven optimization on larger datasets eventually mitigates differences across architectures.